OpenStudy (anonymous):

I forgot how to find the equation of the line in slope-intercept form (3,-1) (-1,5) can you refresh my memory

OpenStudy (anonymous):

use point slope form

OpenStudy (anonymous):

then get y on one side of the equation.

OpenStudy (anonymous):

is that like mx+b sorry I haven't been in school in a while and im relearning it all over again

OpenStudy (anonymous):

it is y-y_1=m(x-x_1)

OpenStudy (anonymous):

this is what it looks like \[(y-y_{1})=m(x-x_{1})\] m is slope y1 is y coordinate x! is xcoordinate

OpenStudy (anonymous):

us your 2 points to find the slope a substitute one of the points into the formula and get y on one side of the equation

OpenStudy (anonymous):

im confussed how would I do that with (3,-1) (-1,5)

OpenStudy (anonymous):

use slope formula substitute into the formula I gave you take one of the points substitute into the formula

OpenStudy (anonymous):

this will help https://www.youtube.com/watch?v=DtPhV8q2D2E

OpenStudy (anonymous):

I think I got it 2y=-3x+7

OpenStudy (anonymous):

This video will solve the process. ignore the first one use this as a guide https://www.youtube.com/watch?v=bSw5YYKZXrA

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I hope it helps. if you still do not understand, try using online learning resources.

jimthompson5910 (jim_thompson5910):

2y=-3x+7 is very close. You just need to get that into y = mx+b form

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

what do you get when you do that?